# Restoring Missing Modes of 21cm Intensity Mapping with Deep Learning: Impact on BAO Reconstruction

### v1

@@ 5 Conclusions and Discussion
With recent advancements in detecting the auto-power spectrum of 21cm intensity mapping [73], we are approaching closer to the ultimate goal of cosmological constraints using this technique. Despite these advances, many aspects of the data analysis process remain unexplored, potentially impacting future measurements. This paper focuses on the BAO reconstruction technique, originally developed for galaxy redshift surveys, which has evolved over many years to effectively reverse the effects of non-linear structure formation and restore the linear BAO signature, thereby enhancing the measurement accuracy of the sound horizon scale. However, foreground contamination severely compromises modes within the Fourier regions affected, including both intrinsic $k_{\parallel}<k_{\rm cut}$ and a wedge-like region, leading to the loss of valuable cosmological information and complicating the application of BAO reconstruction, which fundamentally operates in the configuration space. Fortunately, however, the non-linear structure formation, results in complex mode coupling, enabling the foreground-unaffected modes carry information from those that are missing.
@@
In this paper, we utilize the U-Net deep learning architecture to retrieve this missing information and assess its impact on BAO reconstruction. We demonstrate that this popular and powerful deep learning architecture is capable of retrieving the missing information entangled within the observed data. Despite the success demonstrated here, there is potential for further improving the model by optimizing the loss function, introducing more effective hierarchical structure, and providing more training dataset etc. During our training process, we discovered an interesting scale invariance when applying the AI model to large-scale structure data: a model trained on a coarser map could be successfully applied to finer resolutions, achieving even higher cross-correlation ratios. We attribute this to the scale-invariance in the mode-coupling of the non-linear structure formation process, where the patterns learned by the deep learning model at coarser grid resolutions represent meaningful physical processes applicable across different scales. Furthermore, the architecture of U-Net, with its hierarchical structure of convolutional layers that capture features at multiple scales and then reconstruct the output at higher resolutions, is particularly well-suited to leveraging the scale-invariance of large-scale structure data.
@@
Despite the promising performance demonstrated here, many systematic effects could potential affect the result. Crucially, for a machine learning-based method, it is essential to incorporate real-world observational effects into the training process to mitigate potential biases in the trained model. These include factors such as instrumental beam pattern, thermal noise, and residual foregrounds, among others. We plan to address these aspects in future studies. Meanwhile, this is also the reason why we adopted this two-step approach of reconstruction, rather than directly predicting the linear fields from the foreground-contaminated temperature map. Given physical limitation set by the shell-crossing, the deep learning process is unlikely to significantly outperform traditional BAO reconstruction algorithms. Consequently, using AI in the subsequent process might introduce additional cosmology-dependent errors without significant benefits.
@@ Acknowledgments
This work is supported by the National SKA Program of China (Grants Nos. 2022SKA0110200, 2022SKA0110202 and 2020SKA0110401), the National Science Foundation of China (Grants Nos. 12473006, 12373005), the China Manned Space Project with No. CMS-CSST-2021 (B01, A02, A03). The authors acknowledge the Beijing Super Cloud Center (BSCC, URL: http://www.blsc.cn/) for providing HPC resources that have contributed to the research results reported within this paper.

### v2

@@ 5 Conclusions and Discussion
With recent advancements in detecting the auto-power spectrum of 21cm intensity mapping [77], we are approaching closer to the ultimate goal of cosmological constraints using this technique. Despite these advances, many aspects of the data analysis process remain unexplored, potentially impacting future measurements. This paper focuses on the BAO reconstruction technique, originally developed for galaxy redshift surveys, which has evolved over many years to effectively reverse the effects of non-linear structure formation and restore the linear BAO signature, thereby enhancing the measurement accuracy of the sound horizon scale. However, foreground contamination severely compromises modes within the Fourier regions affected, including both intrinsic $k_{\parallel}<k_{\rm cut}$ and a wedge-like region, leading to the loss of valuable cosmological information and complicating the application of BAO reconstruction, which fundamentally operates in the configuration space. Fortunately, the non-linear structure formation, results in complex mode coupling, enabling the foreground-unaffected modes carry information from those that are missing.
@@
In this paper, we utilize the U-Net deep learning architecture to retrieve this missing information and assess its impact on BAO reconstruction. We demonstrate that this popular and powerful deep learning architecture is capable of retrieving the missing information entangled within the observed data. Despite the success demonstrated here, there is potential for further improving the model by optimizing the loss function, introducing more effective hierarchical structures, and providing more training datasets etc. During our training process, we discovered an interesting scale invariance when applying the AI model to large-scale structure data: a model trained on a coarser map could be successfully applied to finer resolutions, achieving even higher cross-correlation ratios. This indicates a successful restoration of the phase angles of the corresponding Fourier modes. One possible explanation is the scale invariance in mode coupling inherent in the nonlinear structure formation process. The deep learning model may be learning patterns at coarser grid resolutions that represent fundamental physical processes, allowing them to generalize across different scales. Furthermore, the architecture of U-Net, with its hierarchical structure of convolutional layers that capture features at multiple scales and then reconstruct the output at higher resolutions, is particularly well-suited to leveraging the scale-invariance of the large-scale structure data. However, this comes at the cost of introducing additional scale dependence in the power spectrum, as shown in Figure 3. Further investigation is needed to fully understand the underlying cause of this effect.
However, as discussed earlier, our approach has several limitations, particularly in modeling the HI distribution for training samples. First, the mode coupling in COLA simulations may deviate from that in full N-body simulations, potentially affecting the AI model’s performance when applied to real data. Second, the low mass resolution of our dark matter field necessitates subgrid modeling of HI, which fails to preserve the spatial correlations present in a physically realistic HI distribution. This could introduce biases and alter mode coupling, particularly impacting AI restoration in high-resolution maps. Finally, our approximations in modeling redshift-space distortions may introduce additional deviations. In both resolved and unresolved halos, HI velocities deviate from those of their host galaxies or the underlying dark matter field, adding further uncertainties. To improve our approach, future efforts could focus on enhancing the accuracy of both dark matter and HI field modeling. Additionally, given the challenges in generating large, high-precision training datasets, one could conduct consistency tests and apply the AI-trained model to hydrodynamical simulation data for further validation.
@@
Furthermore, there are many systematic effects could potential affect the future application of this approach. Crucially, for a machine learning-based method, it is essential to incorporate real-world observational effects into the training process to mitigate potential biases in the trained model. These include factors such as instrumental beam pattern, thermal noise, and residual foregrounds, among others. We plan to address these aspects in future studies. Meanwhile, this is also the reason why we adopted this two-step approach of reconstruction, rather than directly predicting the linear fields from the foreground-contaminated temperature map. Given physical limitations set by the shell-crossing, the deep learning process is unlikely to significantly outperform traditional BAO reconstruction algorithms. Consequently, using AI in the subsequent process might introduce additional cosmology-dependent errors without significant benefits.
@@ Acknowledgments
This work is supported by the National SKA Program of China (Grants Nos. 2022SKA0110200, 2022SKA0110202 and 2020SKA0110401), the National Science Foundation of China (Grants Nos. 12473006, 12373005, 12261141691), the China Manned Space Project with No. CMS-CSST-2021 (B01, A02, A03). The authors acknowledge the Beijing Super Cloud Center (BSCC, URL: http://www.blsc.cn/) for providing HPC resources that have contributed to the research results reported within this paper.
Appendix A Parametric Fitting of BAO Signature
As discussed in Section 4.2, we followed the standard approach for extracting the BAO damping parameter $\Sigma$ while accounting for the scale-dependent distortions introduced by the AI model. Specifically, Figure 7 shows the multi-dimensional parameter distributions from the MCMC fitting of the AI-restored power spectrum without BAO reconstruction, while Figure 8 presents the corresponding results with both AI restoration and BAO reconstruction.
Figure 7: The distribution of model parameters from the MCMC analysis for the AI-restored power spectrum without BAO reconstruction. The polynomial coefficients $a_{i}$ exhibit mostly Gaussian distributions, whereas the damping parameter $\Sigma$ shows some degree of non-Gaussianity.
Figure 8: Similar to Figure 7, the distribution of model parameters from the MCMC analysis for AI-restored power spectrum with BAO reconstruction.